OpenStudy (lilmike234):

The population in millions of a certain country can be approximated by the function P(x)=40 x 1.02^x Where x is the number of years after 2000. Which of the following calculations will tell in what year the population can be expected to reach 200 million? 1. In(5)/In(1.02) 2. In(5/1.02) 3. In(5/1.02)+2000

OpenStudy (lilmike234):

Or In(5)/In(1.02)+2000

OpenStudy (anonymous):

200 = 40 * 1.02^x 200/40 = 1.02^x 5 = 1.02^x ln(5) = ln( 1.02^x) ln(5)= x * ln(1.02) ln(5)/ ln(1.02) = x Did you get this far?

OpenStudy (lilmike234):

Yeah and then I got stuck after this